OpenStudy (anonymous):

Suppose you got 8 mangoes and 3 apples for $18 and 5 apples and 3mangoes for $14.50 how much does each apple cost and how much does each mango cost?

OpenStudy (anonymous):

setting mangoes = m and apples = a, you have two equations 8m + 3a = 18 5a + 3m = 14.5 then, from the first equation: a = (18-8m)/3 = 6 - (8/3)m Plug it in the second equation: 5(6 - 8m/3) + 3m = 14.5 30 - 40m/3 + 3m = 14.5 31m/3 = 15.5 m = 1.5 plug this back into the original equation 8(1.5) + 3a = 18 a = 2 There you go :)